# Branching Flows: Discrete, Continuous, and Manifold Flow Matching with Splits and Deletions

### v1

@@ D.2 The Branching Infinitesimal Generator
$\mathcal{L}^{\mathrm{branch}}_{t}f(x,n)$ $=\mathcal{L}_{t}^{\mathrm{jump}}f(x,n)+\mathcal{L}_{t}^{\mathrm{base}}f(x,n)$
$=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}(x,n)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}(x,n)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\mathrm{base}}(x,n)\rangle_{t,(x,n)}^{\mathrm{base}}.$
D.3 The Branching Flows Loss
@@
In this section, we simply denote $x_{t}\in\bigsqcup_{n\geq 1}\mathcal{E}$ for a general component-wise state space, augmented or not. The conditional branching flows loss is the conditional generator matching loss associated to the linear parametrization from Appendix D.2. We specify a conditional infinitesimal generator of the form
$\mathcal{L}^{Z,\mathrm{branch}}_{t}f(x,n)$ $=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}^{Z}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}^{Z}(x)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{Z,\mathrm{base}}(x)\rangle_{t,(x,n)}^{\mathrm{base}}.$
@@
$\mathcal{L}^{\theta,\mathrm{branch}}_{t}f(x,n)$ $=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}^{\theta}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}^{\theta}(x)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\theta,\mathrm{base}}(x)\rangle_{t,(x,n)}^{\mathrm{base}}.$
@@
As in LPAL, when a generator is expressed as a sum of linear parametrizations, a valid per-term conditional generator matching loss is then the sum of Bregman divergences
$D^{\mathrm{split}}_{t,n}(R_{t}^{Z}(x,n),R_{t}^{\theta}(x,n))+D^{\mathrm{del}}_{t,n}(\rho_{t}^{Z}(x,n),\rho_{t}^{\theta}(x,n))+D^{\mathrm{base}}_{t,n}(F_{t}^{Z}(x,n),F_{t}^{\theta}(x,n)).$
@@
Assuming that $R_{t}^{Z}(x,n)$ can be interpreted as the remaining number of splits for each component and that $\rho_{t}^{Z}(x,n)$ can be interpreted as deletion probabilities, we have that $R_{t}^{Z}(x,n)\in[0,\infty)^{n}$ and $\rho_{t}^{Z}(x,n)\in[0,1]^{n}$. This allows us to choose a Poisson-like Bregman for splits and a cross-entropy loss for deletions, as it is shown in LPAL. If we moreover assume that the base Bregman divergence is separable and splits along each component, the conditional branching flows loss becomes:
$L_{\mathrm{CBF}}(\theta)=\mathbb{E}_{t\sim\mathrm{\mathcal{D}[0,1]},Z\sim p_{Z},X_{t}\sim p_{t|Z}(dx|z)}\Big[$ $\sum_{i=1}^{N}\Big({R_{t,i}^{\theta}(X_{t})-R_{t,i}^{Z_{c}}(X_{t})\log R_{t,i}^{\theta}(X_{t})}\Big)$
$+\sum_{i=1}^{N}\Big(-[\rho^{Z}_{t,i}(X_{t})\log\rho_{t,i}^{\theta}+(1-\rho^{Z}_{t,i}(X_{t}))\log(1-\rho_{t,i}^{\theta}(X_{t}))]\Big)$
$+\sum_{i=1}^{N}\Big(D_{t,i}^{\mathrm{base}}\big(F_{t,i}^{Z,\mathrm{base}}(X_{t}),F_{t,i}^{\theta,\mathrm{base}}(X_{t})\big)\Big)\Big].$
Appendix E Further Empirical Details
E.1 QM9 analysis
10k samples from the Branching Flows QM9 model were generated with a step schedule $t=1-(\cos(\pi s)+1)/2$ for $s\in(0,1/1000,2/1000,...1)$.
@@ E.1 QM9 analysis
We study the QM9 Branching Flows model outputs by comparison to the QM9 data, comparing marginal distributions of atom counts (total, carbon, nitrogen, oxygen, flourine, and hydrogen), and molecular properties. Atom counts are read directly from the .xyz model output or QM9 data.
@@
We used the model from Campbell2023TransDimensionalGM to generate 10,000 samples, and analyze them as above. Note: the UMAPs, which are computed jointly from all the samples, will differ in the analysis that includes transdimensional jump samples.
To quantify univariate distribution overlap, we use a Kolmogorov-Smirnov-like statistic between the empirical cumulative distribution functions of samples $Y$ from each model against $QM9$ data. We use $1-KS_{D}$, where
@@
E.2 Antibodies
E.2.1 Antibody Sequence Generation: Metrics and Comparisons
To evaluate Branching Flows, 10 00010\,000 sequences were generated with 200 uniformly-spaced steps from Branching Flows and Oracle Length models and then compared to natural ones. Sequence diversity was assessed by computing the minimum pairwise cosine distance (over vectors of 3-mer counts) within each set. Sequence ‘novelty’ was evaluated by computing the minimum Levenshtein distance to any sequence in a hold-out validation data set (needed because otherwise the novelty of all natural sequences would be zero).

### v3

@@ E.2 The Marginal Branching Infinitesimal Generator Rewrite
$\mathcal{L}^{\mathrm{branch}}_{t}f(x)$ $=\mathcal{L}_{t}^{\mathrm{jump}}f(x)+\mathcal{L}_{t}^{\mathrm{base}}f(x)$
$=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}(x)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\mathrm{base}}(x)\rangle_{t,x}^{\mathrm{base}}.$
E.3 Conditioning on Latent Discrete Processes
In Branching Flows, the branching structure is encoded in the trees in $Z$, but the timing of the bifurcation events is stochastic. This means that, given the trees in $Z$ and a state $X_{t}$, each element in $X_{t}$ cannot be unambiguously assigned to a branch on a tree. We thus need to augment the state with a branch-tracking index $G_{t}$. While it is possible to additionally condition the model on the branch tracking index, this places an architectural burden on the model. To avoid this, we learn a generator on $X$ that marginalizes over the time-dependent auxiliary process $G_{t}$, extending the framework of Edit Flows (Havasi2025EditFF) from the purely discrete $X_{t}$ case to where $X_{t}$ can be any process allowed under Generator Matching, including drift, diffusion, and jumps.
With joint conditional paths $(X_{t},G_{t})|(Z=z)\sim p_{t\mid Z}(\mathrm{d}x,\mathrm{d}g\mid z)$, we train a model that marginalizes over $G_{t}$ and $Z$, learning a marginal generator $\mathcal{L}_{t}$ that generates a process $\widehat{X}_{t}$ with marginals $p_{t}(\mathrm{d}x)$ that generates samples from the data distribution $q(dx)$. In what follows, we will assume that the Generator Matching regularity conditions apply to the marginal process $\widehat{X}_{t}$ and the joint marginal $(X_{t},G_{t})$ (marginalizing over $Z$), and the joint conditional $(X_{t},G_{t})|(Z=z)$.
Let $\mathcal{G}$ be discrete with $|\mathcal{G}|<\infty$ and let $(X_{t},G_{t})$ be an $S\times\mathcal{G}$-valued process generated by $W_{t}:\mathcal{T}(S\times\mathcal{G})\to C(S\times\mathcal{G},\mathbb{R})$. For fixed $g\in\mathcal{G}$, define $\mathcal{L}_{t}^{g}f(x):=W_{t}\tilde{f}(x,g)$, where $\tilde{f}(x,g)=f(x)$.
In what follows, we define the measure $\mu^{g}_{X_{t}}(A):=p_{X_{t},G_{t}}(A\times\{g\})$. Note that for each $E\subset\mathcal{G}$ we have $p_{X_{t},G_{t}}(dx,E)=\sum_{g\in E}\mu^{g}_{X_{t}}(dx)$. The following lemma holds:
Lemma E.3.1.
$\partial_{t}\sum_{g\in\mathcal{G}}\int_{S}f(x)\mu^{g}_{X_{t}}(dx)=\sum_{g\in\mathcal{G}}\int_{S}\mathcal{L}_{t}^{g}f(x)\mu^{g}_{X_{t}}(dx).$
Proof.
Write
$\partial_{t}\sum_{g\in\mathcal{G}}\int_{S}f(x)\mu^{g}_{X_{t}}(dx)$ $=\partial_{t}\int_{S}f(x)\sum_{g\in\mathcal{G}}\mu^{g}_{X_{t}}(dx)$
where in the above we are integrating with respect to the measure $\sum_{g\in\mathcal{G}}\mu_{X_{t}}^{g}(dx)$. This becomes
$=\partial_{t}\int_{S}f(x)p_{t}(dx,\mathcal{G})$
$=\partial_{t}\int_{S}f(x)\int_{\mathcal{G}}p_{t}(dx,dg)$
$=\partial_{t}\int_{S\times\mathcal{G}}f(x)p_{t}(dx,dg)$
$=\partial_{t}\int_{S\times\mathcal{G}}\tilde{f}(x,g)p_{t}(dx,dg)$
and since $W_{t}$ generates $p_{t}(dx,dg)$, by the Kolmogorov Forward Equation, it holds:
$=\int_{S\times\mathcal{G}}W_{t}\tilde{f}(x,g)p_{X_{t},G_{t}}(dx,dg)$
$=\sum_{g\in\mathcal{G}}\int_{S}\mathcal{L}_{t}^{g}f(x)\mu^{g}_{X_{t}}(dx).$
Theorem E.3.2.
It holds that
$\mathcal{L}_{t}f(x):=\sum_{g\in\mathcal{G}}\mathcal{L}_{t}^{g}f(x)\,p_{G_{t}|X_{t}}(g|x)$
generates $p_{t}(dx)$, i.e., $\mathcal{L}_{t}$ solves the KFE for $p_{t}$:
$\partial_{t}\int f(x)p_{t}(dx)=\int\mathcal{L}_{t}f(x)p_{t}(dx).$
Moreover, assume that $F_{t}^{g}:S\to\Omega\subset V$ is a linear parametrization of $\mathcal{L}_{t}^{g}f(x)$ over a vector space $V$. If there is a linear $\mathcal{K}:\mathcal{T}\to\mathcal{C}(S;V)$ on the inner product space $\langle\cdot,\cdot\rangle_{V}$ such that
$\mathcal{L}_{t}^{g}f(x)=\langle\mathcal{K}f(x),F_{t}^{g}(x)\rangle_{V}$
for each $g\in\mathcal{G}$, then
$F_{t}(x)=\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}[F_{t}^{G_{t}}(x)]$
linearly parametrizes $\mathcal{L}_{t}$.
Proof.
For each $t\in[0,1]$, we have $p_{t}(dx)=\sum_{g\in\mathcal{G}}\mu^{g}_{X_{t}}(dx)$, and hence:
$\partial_{t}\int_{S}f(x)p_{t}(dx)$ $=\partial_{t}\int_{S}f(x)\sum_{g\in\mathcal{G}}\mu^{g}_{X_{t}}(dx)$
$=\partial_{t}\sum_{g\in\mathcal{G}}\int_{S}f(x)\mu^{g}_{X_{t}}(dx)$
and by Lemma E.3.1, it holds:
$=\sum_{g\in\mathcal{G}}\int_{S}\mathcal{L}_{t}^{g}f(x)\mu^{g}_{X_{t}}(dx)$
$=\sum_{g\in\mathcal{G}}\int_{S}\mathcal{L}_{t}^{g}f(x)p_{G_{t}|X_{t}}(g|x)p_{t}(dx)$
$=\int_{S}\underbrace{\left(\sum_{g\in\mathcal{G}}\mathcal{L}_{t}^{g}f(x)p_{G_{t}|X_{t}}(g|x)\right)}_{=\mathcal{L}_{t}f(x)}p_{t}(dx).$
By the linearity of expectations and the map $w\mapsto\langle v,w\rangle_{V}$ for fixed $v\in V$, we have
$\mathcal{L}_{t}f(x)$ $=\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}[\mathcal{L}_{t}^{G_{t}}f(x)]$
$=\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}[\langle\mathcal{K}f(x),F_{t}^{G_{t}}(x)\rangle_{V}]$
$=\langle\mathcal{K}f(x),\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}[F_{t}^{G_{t}}(x)]\rangle_{V}$
$=\langle\mathcal{K}f(x),F_{t}(x)\rangle_{V}.$
Remark E.3.3.
As shown above, the process $\widehat{X}_{t}$ generated by $\mathcal{L}_{t}$ has marginals $p_{t}(dx)$, but we simply write our expectations over $X_{t}\sim p_{t}(dx)$ for clarity of notation.
Theorem E.3.4.
The auxiliary-process-conditioned generator matching loss and the generator matching loss coincide up to a constant in $\theta$:
$\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D},X_{t},G_{t}\sim p_{t}(dx,dg)}D(F_{t}^{G_{t}}(X_{t}),F_{t}^{\theta}(X_{t}))=\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D},X_{t}\sim p_{t}(dx)}[D(F_{t}(X_{t}),F_{t}^{\theta}(X_{t}))].$
Proof.
Since expectations commute with the left slot of Bregman divergences under the gradient (cf. Lipman2024FlowMG, billera2025timedependentlossreweighting), we obtain:
$\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D},X_{t},G_{t}\sim p_{t}(dx,dg)}D(F_{t}^{G_{t}}(X_{t}),F_{t}^{\theta}(X_{t}))$ $=\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D},X_{t}\sim p_{t}(dx)}\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}D(F_{t}^{G_{t}}(X_{t}),F_{t}^{\theta}(X_{t}))$
$=\mathbb{E}_{t\sim\mathcal{D},X_{t}\sim p_{t}(dx)}\nabla_{\theta}\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}D(F_{t}^{G_{t}}(X_{t}),F_{t}^{\theta}(X_{t}))$
$=\mathbb{E}_{t\sim\mathcal{D},X_{t}\sim p_{t}(dx)}\nabla_{\theta}D(\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t}}(dg|x)}[F_{t}^{G_{t}}(X_{t})],F_{t}^{\theta}(X_{t}))$
$=\mathbb{E}_{t\sim\mathcal{D},X_{t}\sim p_{t}(dx)}\nabla_{\theta}D(F_{t}(X_{t}),F_{t}^{\theta}(X_{t}))$
$=\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D},X_{t}\sim p_{t}(dx)}D(F_{t}(X_{t}),F_{t}^{\theta}(X_{t})).$
Remark E.3.5.
The arguments in Holderrieth2024GeneratorMG allow us to include an additional latent conditioning variable $z\in\mathcal{Z}$ that is often used to steer the path towards terminating at a particular point of the data distribution. Let $\mathcal{L}_{t}^{z,g}$ denote the conditional generator given $z\in\mathcal{Z}$ and $g\in\mathcal{G}$. Define the $z$-conditioned generator
$\mathcal{L}_{t}^{z}f(x):=\mathbb{E}_{G_{t}\sim p_{G_{t}|X_{t},Z}(dg|x,z)}[\mathcal{L}_{t}^{z,G_{t}}f(x)].$
Then, by Theorem E.3.2 and Proposition 1 of GM,
$\mathbb{E}_{Z\sim p_{Z|X_{t}}(dz|x),G_{t}\sim p_{G_{t}|Z,X_{t}}(dg|z,x)}[\mathcal{L}_{t}^{Z,G_{t}}f(x)]$ $=\mathbb{E}_{Z\sim p_{Z|X_{t}}(dz|x)}[\mathbb{E}_{G_{t}\sim p_{G_{t}|Z,X_{t}}(dg|z,x)}[\mathcal{L}_{t}^{Z,G_{t}}f(x)]]$
$=\mathbb{E}_{Z\sim p_{Z|X_{t}}(dz|x)}[\mathcal{L}_{t}^{Z}f(x)]$
$=\mathcal{L}_{t}f(x),$
so averaging over a fixed $z\in\mathcal{Z}$ and a discrete process $(G_{t})_{0\leq t\leq 1}$ recovers the same marginal generator $\mathcal{L}_{t}$. For the loss, if we write $F_{t}^{z,g}$ for a linear parametrization of $\mathcal{L}_{t}^{z,g}$, the same linearity argument as in Theorem E.3.4 gives
$\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D}[0,1],X_{t},Z,G_{t}}D(F_{t}^{Z,G_{t}}(X_{t}),F_{t}^{\theta}(X_{t}))$ $=\mathbb{E}_{t\sim\mathcal{D},X_{t}}[\nabla_{\theta}\mathbb{E}_{Z\sim p_{Z|X_{t}}(dz|x),G_{t}\sim p_{G_{t}|Z,X_{t}}(dg|z,x)}[D(F_{t}^{Z,G_{t}}(X_{t}),F_{t}^{\theta}(X_{t}))]]$
$=\mathbb{E}_{t\sim\mathcal{D},X_{t}}[\nabla_{\theta}D(\mathbb{E}_{Z\sim p_{Z|X_{t}}(dz|x),G_{t}\sim p_{G_{t}|Z,X_{t}}(dg|z,x)}[F_{t}^{Z,G_{t}}(X_{t})],F_{t}^{\theta}(X_{t}))]$
$=\mathbb{E}_{t\sim\mathcal{D},X_{t}}\nabla_{\theta}D(F_{t}(X_{t}),F_{t}^{\theta}(X_{t}))$
$=\nabla_{\theta}\mathbb{E}_{t\sim\mathcal{D},X_{t}}D(F_{t}(X_{t}),F_{t}^{\theta}(X_{t})),$
so conditioning on a fixed $z\in\mathcal{Z}$ does not change the gradient of the loss.
E.4 The Branching Flows Loss
@@
The conditional branching flows loss is the auxiliary-process-conditioned generator matching loss associated to the linear parametrization from Appendix E.2, with conditional dynamics as described in Appendix E.3. For $x\in\bigsqcup_{n\geq 1}\mathcal{E}^{n}$, we specify a conditional infinitesimal generator of the form
$\mathcal{L}^{z,g,\mathrm{branch}}_{t}f(x)$ $=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}^{z,g}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}^{z,g}(x)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{z,g,\mathrm{base}}(x)\rangle_{t,x}^{\mathrm{base}}.$
@@
$\mathcal{L}^{\theta,\mathrm{branch}}_{t}f(x)$ $=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}^{\theta}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}^{\theta}(x)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\theta,\mathrm{base}}(x)\rangle_{t,x}^{\mathrm{base}}.$
@@
As in (billera2025timedependentlossreweighting), when a generator is expressed as a sum of linear parametrizations, a valid per-term conditional generator matching loss is then the sum of Bregman divergences
$D^{\mathrm{split}}_{t,x}(R_{t}^{z,g}(x),R_{t}^{\theta}(x))+D^{\mathrm{del}}_{t,x}(\rho_{t}^{z,g}(x),\rho_{t}^{\theta}(x))+D^{\mathrm{base}}_{t,x}(F_{t}^{z,g,\mathrm{base}}(x),F_{t}^{\theta,\mathrm{base}}(x)),$
@@
again noting that $x$ stands for $x\in S$. Assuming that $R_{t}^{z,g}(x)$ can be interpreted as the remaining number of splits ahead of each element and that $\rho_{t}^{z,g}(x)$ can be interpreted as deletion probabilities, we choose a Poisson-like Bregman for splits and a cross-entropy loss for deletions.
Note that these loss functions require a notion of Bregman divergence associated with convex functions that are only differentiable on a subset of their domain, as considered in (billera2025timedependentlossreweighting)
If we also assume that the base Bregman divergence is separable and splits along each ahead, the conditional branching flows loss becomes:
$L_{\mathrm{CBF}}(\theta)$ $=\mathbb{E}_{t\sim\mathcal{D}[0,1],Z\sim p_{Z},(X_{t},G_{t})\sim p_{X_{t},G_{t}|Z}(dx,dg|z)}$
$\qquad\Big[\sum_{i=1}^{L_{t}}\Big({R_{t,i}^{\theta}(X_{t})-R_{t,i}^{Z,G_{t}}(X_{t})\log R_{t,i}^{\theta}(X_{t})}\Big)$
$\qquad+\sum_{i=1}^{L_{t}}\Big(-[\rho^{Z,G_{t}}_{t,i}(X_{t})\log\rho_{t,i}^{\theta}(X_{t})+(1-\rho^{Z,G_{t}}_{t,i}(X_{t}))\log(1-\rho_{t,i}^{\theta}(X_{t}))]\Big)$
$\qquad+\sum_{i=1}^{L_{t}}\Big(D_{t,X_{t},i}^{\mathrm{base}}(F_{t,i}^{Z,G_{t},\mathrm{base}}(X_{t}),F_{t,i}^{\theta,\mathrm{base}}(X_{t}))\Big)\Big].$
Appendix F Empirical Details and Supplementary Results
F.1 QM9 analysis
F.1.1 Data
F.1.2 Branching Flows specification
  • Conditional base process:
    • Continuous: OU process flow (Appendix D.1), with a mean-reversion rate $\theta=5$, and where the variance decays from 10 when $t=0$ to $0.001$ by $t=1$.
    • Discrete: DFM convex interpolation (appendix D.2) of $x_{0}$, uniform noise, and $x_{1}$, with hazards $\mathcal{D}_{1}=\text{Beta}(2,2),\mathcal{D}_{2}=\text{Beta}(2,2),\omega_{u}=0.2$.
  • Branching hazard distribution: $\operatorname{Beta}(1,3/2)$
  • Deletion hazard distribution: $U(0,1)$
  • $x_{0}$ length distribution: 1, always.
F.1.3 Model and Training
We use a 12 layer transformer model (12 heads, each with head dimension of 64), with an embedding dimension of 384. The atom positions are encoded by Random Fourier Features (RFFs), and pairwise spatial features are also provided to each layer by a learnable, layer-specific, head-specific attention bias. Rotary Positional Encoding (RoPE, su2023roformerenhancedtransformerrotary) is used to encode the primary sequence ordering. The final six layers additively update the atom positions (and the spatial pair features are recalculated after each position update). The model outputs the endpoint-predicted (i.e. by $t=1$) locations, logits over the endpoint-predicted discrete state, the log of the per-element expected number of future bifurcations by $t=1$, and a deletion binary logit. The model was trained, using Muon (jordan2024muon) with post-warmup learning rate of $0.005$, for 500k batches (where each batch was 128 randomly-sampled QM9 molecules) and a linear learning-rate cooldown for the last 50k batches. For sampling, 10k samples from the Branching Flows QM9 model were generated with a time step schedule $t=1-(\cos(\pi s)+1)/2$ for $s\in(0,1/1000,2/1000,\ldots,1)$.
@@ F.1.3 Model and Training
We study the QM9 Branching Flows model outputs by comparison to the QM9 data, comparing distributions of atom counts (total, carbon, nitrogen, oxygen, fluorine, and hydrogen), and molecular properties. Atom counts are read directly from the .xyz model output or QM9 data.
@@
Additionally, we used the model from Campbell2023TransDimensionalGM to generate 10,000 samples, and analyze them as above. Note: the UMAPs, which are computed jointly from all the samples, will differ in the analysis that includes transdimensional jump samples. This is compared to Branching Flows in table 1 and figure 9.
To quantify distribution overlap, we use a Kolmogorov-Smirnov-like statistic between the empirical cumulative distribution functions of samples $Y$ from each model against QM9 data. We use $1-KS_{D}$, where
@@
F.2 Antibodies
F.2.1 Data
F.2.2 Branching Flows Specification
  • Conditional base process:
    • Discrete: DFM convex interpolation (appendix D.2) of $x_{0}$, uniform noise, and $x_{1}$, with $\omega_{u}=0.2$, and CDFs $F_{1}$ and $F_{2}$ of $\mathcal{D}_{1}=\text{Beta}(1.5,1.5),$ and $\mathcal{D}_{2}=\text{Beta}(2,2),$ respectively.
  • Branching hazard distribution: $\operatorname{Beta}(1,2)$
  • Deletion hazard distribution: $U(0,1)$
  • To-be-deleted policy: deletion rate $d_{r}=1.2$
  • $x_{0}$ state distribution: masked tokens.
  • $x_{0}$ length distribution: uniform between 110 and 140.
F.2.3 Model and Training
The models (Branching Flows, oracle-length and Edit Flows) were trained using a batch size of 64 resulting in a total of 234 375234\,375 training steps for the 15 million sequence dataset. Muon was used with a 2000 step linear warm-up followed by a cosine decay starting at $10^{-3}$ and decaying to $10^{-7}$.
F.2.4 Antibody Sequence Generation: Metrics and Comparisons
To evaluate Branching Flows, 10 00010\,000 sequences were generated with 1 0001\,000 uniformly-spaced steps from Branching Flows, Oracle Length, and Edit Flows models and then compared to natural ones. Sequence diversity was assessed by computing the minimum pairwise cosine distance (over vectors of 3-mer counts) within each set. Sequence ‘novelty’ was evaluated by computing the minimum Levenshtein distance to any sequence in a hold-out validation data set (needed because otherwise the novelty of all natural sequences would be zero).